OpenStudy (anonymous):

3l+y=5+5l solve for l. More insight on that, please?

OpenStudy (anonymous):

l. this means the absolute value of the number as in 3 is the absolute value +y=5+5 as in 3+y=5+5 simply just simplify those i do think thats it unless thats a L

OpenStudy (anonymous):

sorry i should've clarify that it's a literal equation the "l" are L's

OpenStudy (jdoe0001):

I'm thinking is a letter "l"

OpenStudy (anonymous):

oh okay haha thanks well so its 3L+y=5+5L all you need to do is get the same variables on the same side as in 3L-3L+y=5+5L-3L which is y=2L+5

OpenStudy (anonymous):

you have to isolate the value make y alone do you get what im trying to say ?

OpenStudy (anonymous):

*variable

OpenStudy (anonymous):

then what else you do is you want to get L alone right ?

OpenStudy (anonymous):

y=2L+5 what do you think you do next ?

OpenStudy (anonymous):

yea but is was to solve for L. I got L=5-y/-2

OpenStudy (anonymous):

\[l=\frac{ 5-y }{ -2 }\]

OpenStudy (anonymous):

good now you have your answer

OpenStudy (anonymous):

but wait where did you get -2 ?

OpenStudy (anonymous):

i meant over 3

OpenStudy (anonymous):

over -3

OpenStudy (anonymous):

no you were correct the first time as 2 because if you dont remember what we did before was we subtracted 5L-3L which is 2 so the answer would end up being L=5-y/2

OpenStudy (anonymous):

not a negative

OpenStudy (anonymous):

do you get it now ?

OpenStudy (anonymous):

thanks' i just needed the clarification of that negative. Thanks!!

OpenStudy (anonymous):

your welcome anytime